OpenStudy (anonymous):

Please guys help me!! Even if u know only one point ! Given the equation kx+3y−2=0 (a) Find the value k such that the point (−1, 5) belongs to the line (b) Find the value of p such the previous line is perpendicular to 3x − py + 7 = 0. (c) Find the points of intersection.

jhonyy9 (jhonyy9):

a) (-1,5) mean that for x= -1 will get y= -5 so than kx+3y=2 3y=2-kx y=(2-kx)/3 -5 = (2-k(-1))/3 -5 = (2+k)/3 -15= 2+k -17 = k hope these will help you